# MFIT : Multi-FIdelity Thermal Modeling for 2.5D and 3D Multi-Chiplet Architectures

### v2

@@ V-B2 HotSpot [15] configuration:
We also compare our proposed models to the state-of-the-art tool HotSpot [15] for all the system sizes and workload configurations. Since HotSpot was originally designed for 2D chips, we utilize an extension that adds 3D modeling capabilities [53] to model both 2.5D and 3D integrated systems. Geometry and material parameters are set to be identical to our reference FEM model. Since HotSpot does not support thermal conductivity variations in x-y-z directions, we use the average conductivity for anisotropic material layers (ex. C4 bump layer) in the chiplet package. HotSpot also lacks the support for varying grid sizes across different layers. Therefore, we maintain a uniform grid size matching our chiplet layer.
V-C Execution Time Evaluation
@@
2.5D Evaluation: Abstracted FEM simulations take 2.4, 14.5, and 38.0 hours for 16, 36, and 64 chiplet systems, respectively. While providing an accurate reference, these long simulation times and significant development effort make FEM impractical for DSE. Our thermal RC models fill this gap with execution times ranging from 53.0 to 1.8 seconds for 16, 36, and 64 chiplets, as summarized in the first 3 systems in Figure 8. Coupled with the accuracy presented in the previous subsection, the 46 to 136-fold speedup demonstrates their viability as a DSE tool.
@@
Thermal RC models are derived directly from the underlying geometry and material parameters, meaning they can be reconfigured for different hardware and design configurations without re-calibration from FEM simulations. Relaxing this physical system to model connection, our DSS models reduce execution time to 39, 96, and 944 ms for 16, 36, and 64 chiplets respectively. This speedup with respect to the RC model enables runtime temperature prediction which can inform dynamic thermal power management (DTPM) decisions to increase system performance and reliability [54]. While maintaining the accuracy for a given configuration, DSS models need to be regenerated if the sampling period or hardware configuration changes.
3D Evaluation: The execution time results of the 3D system are summarized in the far right system of Figure 8. FEM simulations take approximately 3.3 hours for the single 3D system. The thermal RC model dramatically decreases this runtime to 6 seconds. Similar to the results seen for the 2.5D system, the 3D DSS model again shows a dramatically reduced runtime of 0.07 seconds. For comparison, the execution time of a similarly sized 2.5D system DSS model is 0.09 seconds.
@@
Execution time comparison to HotSpot [15]: The commonly used thermal modeling tool HotSpot belongs to the same class as our thermal RC models. The execution times of the proposed models are significantly faster (1862$\times$ for 16, 607$\times$ for 36, and 245$\times$ for 64 - 2.5D integrated chiplet systems while 817$\times$ for the 16x3 - 3D integrated chiplet system) than HotSpot, as shown in Figure 8.
The significant speedup mentioned above with similar or higher accuracy is primarily attributed to two factors. First, we employ a non-uniform grid for different layers, as explained in Section IV-C. The speedup due to the use of a non-uniform grid size is 13$\times$, 17$\times$, 19$\times$ and 7$\times$ faster execution for 16, 36, 64 - 2.5D integrated and 16x3 - 3D integrated chiplet systems with respect to HotSpot. Second, we employ the adaptive solver LSODA. This solver requires fewer iterations per time step for convergence in comparison to HotSpot. Utilizing an identical grid configuration to HotSpot, the use of this solver alone leads to speedups of approximately 144$\times$, 35$\times$, 13$\times$, and 122$\times$, again with respect to the baseline HotSpot model execution time. Finally, we emphasize that MFIT, our multi-fidelity thermal model set covers a much wider range of accuracy and execution time trade-offs than a specific point solution such as HotSpot.
Figure 9: A heat map of the interposer layer generated by thermal RC model.
V-D Validation of Thermal RC and DSS model
@@
Figure 10: Temperature vs. Time of representative chiplets from 2.5D and 3D systems. WL1 is shown for all systems.
We validate the accuracy of our thermal RC and DSS models by comparing their temperature estimates to full-system FEM simulation results for the same workload and system configurations. This comparison is completed for each of the three 2.5D system sizes and the single 3D system. In addition to the visualization of the temperature estimate over time, two metrics are used to quantify the accuracy of our thermal RC and DSS models against the FEM results. The MAE metric measures the mean absolute error in temperature across the entire simulation duration. Predicting temperature violations (e.g., tracking the time steps when the temperature exceeds the allowed threshold) is often used by DTPM algorithms. Therefore, the second metric measures the accuracy of our models in predicting temperature violations. We set 85C as the maximum allowable temperature threshold for each system without loss of generality [55]. This metric first identifies the time steps in FEM simulations where temperature violations occur (temperature exceeds 85C). Then, it computes the percentage of these violations captured by the thermal RC and DSS models (e.g., 100% means all violations are detected with perfect accuracy). The proposed models conservatively flag violations within one degree of the above-mentioned threshold temperature.
To assist users in visualizing the thermal behavior of the system under test, the RC model also creates a heat map of each layer in the system. As an example, the heat map of the interposer layer of a 2.5D 64 chiplet system is shown in Figure 9. This figure shows the temperature gradient that occurs between the center of the interposer, where the heat producing chiplets are located, and the edges of the system, where there are no chiplets. These maps allow for quick visual verification of the system behavior instead of relying purely on numerical results.
2.5D Validation Results: Figures 10(a), 10(b), and 10(c) plot the temperature as a function of time for each 2.5D system size of a representative chiplet while running workload WL1. All three plots of 2.5D systems clearly show that the systematically constructed thermal RC and DSS models produce near-identical results to the FEM baseline. They closely follow the FEM results during the stress test (temperature increases until reaching the maximum point), randomly changing chiplet power consumption (middle portion), and cool-down periods.
TABLE VII: Mean Absolute Error and accuracy in predicting temperature violations are shown as compared to reference FEM results. The thermal RC, DSS, and HotSpot [15] models are shown.
@@
Accuracy (%)
WL1 Thermal RC 1.23 93.4 1.42 96.9 1.17 99.2 0.82 99.7
DSS 1.23 93.4 1.42 96.9 1.17 99.2 0.82 99.7
HotSpot 2.74 67.7 1.64 96.9 1.69 98.1 1.32 94.5
WL2 Thermal RC 0.86 95.9 1.16 100 1.08 100 1.08 100
DSS 0.86 95.9 1.16 100 1.08 100 1.08 100
HotSpot 1.57 75.0 1.24 100 1.19 100 1.19 100
WL3 Thermal RC 1.02 100 1.28 77.2 1.17 89.3 1.05 100
DSS 1.02 100 1.28 77.2 1.17 89.3 1.05 100
HotSpot 1.97 100 1.18 42.3 1.33 74.4 1.08 100
WL4 Thermal RC 1.41 96.6 1.63 95.5 1.55 97.6 1.30 99.3
DSS 1.41 96.6 1.63 95.5 1.55 97.6 1.30 99.3
HotSpot 2.29 95.6 1.89 92.0 2.15 97.7 1.85 98.0
WL5 Thermal RC 1.01 100 1.25 87.8 1.07 82.4 1.03 100
DSS 1.01 100 1.25 87.8 1.07 82.4 1.03 100
HotSpot 1.94 100 1.16 60.7 1.40 52.9 1.05 100
WL6 Thermal RC 0.89 98.1 1.30 84.8 1.21 90.9 1.11 98.3
DSS 0.89 98.1 1.30 84.8 1.21 90.9 1.11 98.3
HotSpot 1.62 85.8 1.28 89.7 1.56 89.5 1.24 86.7
@@
The first 3 columns of Table VII summarize the accuracy results for all 2.5D systems and workloads. The worst-case mean absolute errors are only 1.41, 1.63, and 1.55 degrees (highlighted in dark red) for 16, 36, and 64 chiplet systems, respectively. These results indicate that the proposed models achieve excellent accuracy across different hardware configurations and workloads. Our models also achieve high accuracy in predicting temperature violations. For example, the worst-case accuracy for the 16-chiplet system is 93.4% (i.e., only 6.6% of the time steps where violations are missed) for WL1. Our models capture 100% of the violations during WL3 and WL6 while missing a handful of them for other workloads. The corresponding accuracy for the 36-chiplet system is above 95% for WL1, WL2, and WL4. We observe 77.2%, 84.8%, and 87.8% accuracy for WL3, WL6, and WL5, respectively. These relatively lower accuracy values stem from sudden temperature spikes that lead to short-term temperature violations in these workload-system combinations. Temperature spikes in specific chiplets occur when several grouped chiplets experience a transient power spike simultaneously. This increases the temperature of the more central chiplets in the group. In these cases, the peak temperatures are mostly at or below 85 degrees with infrequent short-term violations, which can be tolerated. For example, FEM simulations indicate only 255 temperature violations across all chiplets in WL3 compared to 11 thousand violations in WL1. Hence, missing even a few violations impacts the accuracy heavily when the RC and DSS models do not capture these spikes. In contrast, our thermal RC and DSS models effectively detect more prolonged violations, as evidenced by WL1, WL3, and WL4. Similarly, our thermal RC and DSS models achieve high accuracy for the 64-chiplet system, as shown in Table VII. The lowest accuracies are 82.4% for WL5 and 89.3% for WL3, which have few total violations.
@@
3D Validation Results: Figure 10(d) plots the temperature as a function of time for the single 3D system of a representative chiplet running WL1. The plot shows similar behavior to the 2.5D comparison, where the RC and DSS match each other exactly and match the reference FEM results extremely closely during the stress test, random, and cool-down portions of the workload.
@@
The far right column of Table VII summarizes the accuracy results for the single 3D system for each workload. The worst-case MAE is only 1.3 degrees (highlighted again in dark red). This result shows that the proposed models maintain high levels of accuracy even when applied to a stacked die package. This high degree of accuracy is repeated when predicting temperature violations. The worst case temperature violation prediction accuracy is only 98.3%, occurring for workload 6.
Accuracy comparison to HotSpot [15]: Table VII also lists the MAE of HotSpot simulations compared to FEM simulations for 2.5D and the 3D system. For the 2.5D system results, the average error of HotSpot is 0.95, 0.05, and 0.35 degrees greater than our thermal RC and DSS models for the 16, 36, and 64 chiplet systems, respectively. Notably, HotSpot also has lower accuracy in detecting temperature violations, but our primary advantage is in execution time, as discussed in Section V-C. For the 3D results, HotSpot’s average error is 0.22 degrees greater than that of our thermal RC and DSS models. HotSpot again shows lower accuracy in detecting temperature violations in workloads 1, 4, and 6. The larger error across system configurations can be attributed to the tuning of thermal RC model parameters based on the available reference FEM thermal model results. Additionally, HotSpot does not include the ability to provide different thermal conductivities along the x, y, and z-axis, further decreasing accuracy.
VI Conclusion

### v4

@@ 5.2.2. Configuration of other thermal simulators:
We compare our proposed thermal RC and DSS models against the state-of-the-art analytical thermal modeling tools HotSpot (Skadron and otehrs, 2004), 3D-ICE (Huang et al., 2024), and PACT (Yuan et al., 2021) for all system sizes and workload configurations. Since HotSpot was originally designed for 2D chips, we utilize an extension that incorporates 3D modeling capabilities (Meng et al., 2012) to support both 2.5D and 3D integrated systems. Identical geometry and material parameters are set across all tools, aligning them with our reference FEM model. None of these tools support thermal conductivity variations along the x, y, and z directions. Consequently, for anisotropic material layers (e.g., the C4 bump layer) in the chiplet package, we use an averaged thermal conductivity to approximate this configuration. Additionally, HotSpot and PACT do not allow non-uniform grid sizes across different horizontal layers, so a uniform grid size matching our chiplet layer is used. We configure the non-uniform grid in 3D-ICE (Huang et al., 2024) to be consistent with our thermal RC model, ensuring the same number of thermal nodes for each system. Additionally, PACT provides flexibility in solver selection by supporting various SPICE solvers from Xyce (Hutchinson et al., 2002). For consistency, we use the TRAP solver, as given in (Yuan et al., 2021). PACT also enables parallel execution, which we leverage by running all simulations with eight cores on an Intel 10700k CPU.
5.3. Execution Time Evaluation
@@
Figure 8. Execution times (in log scale) of the proposed thermal models and HotSpot (Skadron and otehrs, 2004), PACT (Yuan et al., 2021), and 3D-ICE (Sridhar et al., 2013) for 16, 36, 64 - 2.5D integrated and 16x3 - 3D integrated chiplet systems.
@@
2.5D Evaluation: Abstracted FEM simulations take 2.4, 14.5, and 38.0 hours for 16, 36, and 64 chiplet systems, respectively. While providing an accurate reference, these long simulation times and significant development effort make FEM impractical for DSE. Our thermal RC models fill this gap with execution times ranging from 0.85 to 3.57 seconds for 16, 36, and 64 chiplets, as summarized in the first 3 systems in Fig. 8. Coupled with the accuracy presented in the Section 5.4, the 10,000 to 40,000-fold speedup demonstrates their viability as a DSE tool.
@@
Thermal RC models are derived directly from the underlying geometry and material parameters, meaning they can be reconfigured for different hardware and design configurations without re-calibration from FEM simulations. Relaxing this physical system to model connection, our DSS models reduce execution time to 18, 26, and 54 milliseconds for 16, 36, and 64 chiplets respectively. This speedup with respect to the RC model enables runtime temperature prediction, which can inform dynamic thermal power management (DTPM) decisions to increase system performance and reliability (Brooks et al., 2007). Although accurate real-time power estimation is a critical component of such DTPM frameworks, MFIT treats power as an external input and can be used with power traces obtained from simulators, profilers, or real hardware counters. For instance, physical current sensors that measure runtime power consumption at cluster or core granularities can provide direct input to MFIT (Cochet et al., 2016; Chan and Mok, 2013). Another option is software-based power meters that estimate runtime power by converting performance counter readings into power values, such as Qualcomm’s profiler (qua, 2025). While maintaining the accuracy for a given configuration, DSS can be regenerated from the RC model in a few milliseconds if the sampling period or hardware configuration changes.
3D Evaluation: The execution time results of the 3D system are summarized in the 3D - 16x3 chiplet system of Fig. 8. FEM simulations take approximately 3.3 hours for the single 3D system. The thermal RC model significantly reduces the runtime to 1.6 seconds. Similar to the 2.5D system, the 3D DSS model further reduces runtime to just 24 milliseconds. This substantial speedup is due to the DSS model relying solely on matrix multiplication operations, which are highly optimized on modern hardware through SIMD instructions and other acceleration techniques, such as efficient memory access.
@@
Execution time comparison to existing tools: The commonly used thermal modeling tools, HotSpot (Skadron et al., 2004), 3D-ICE (Huang et al., 2024), and PACT (Yuan et al., 2021) belong to the same class of model as our thermal RC models. As shown in Fig. 8, our proposed thermal RC and DSS methods achieve significantly faster execution times compared to these tools.
HotSpot and PACT lack support for non-uniform grid structures, resulting in more complex ODE formulations and increased computational cost. In particular, HotSpot relies on the classical Runge-Kutta 4 (RK4) method, which is known to be less efficient for stiff systems due to its small time step requirements to maintain stability and accuracy, leading to increased computation time. As a result, our thermal RC model achieves a 4235$\times$ speedup for 16 chiplets, 4082$\times$ for 36 chiplets, 3630$\times$ for 64 chiplets, and 4050$\times$ for the 16$\times$3 3D chiplet system compared to HotSpot. Against PACT, our thermal RC model is 132$\times$ faster for 16 chiplets, 205$\times$ for 36 chiplets, 164$\times$ for 64 chiplets, and 167$\times$ for the 16$\times$3 3D chiplet system.
While 3D-ICE does support non-uniform grids and employs an implicit Euler method for solving transient heat equations, MFIT further improves efficiency by integrating BLAS-optimized routines within the ODE solver pipeline. These routines accelerate matrix-vector multiplications and factorization steps, enabling faster convergence when used alongside the SuperLU solver. In contrast, 3D-ICE relies solely on SuperLU without these enhancements. Consequently, our approach achieves a 3$\times$ speedup for 16 chiplets, 2.3$\times$ for 36 chiplets, 1.9$\times$ for 64 chiplets, and 2.7$\times$ for the 16$\times$3 3D chiplet system over 3D-ICE. Moreover, our DSS method surpasses all RC-based models in execution efficiency, demonstrating the best performance among all evaluated tools. Finally, we emphasize that MFIT, our multi-fidelity thermal model set, covers a much wider range of accuracy and execution time trade-offs than a specific point solution, providing greater flexibility for different modeling needs.
5.4. Validation of Thermal RC and DSS model
@@
Figure 9. Temperatures of representative chiplets from 2.5D and 3D systems while running WL1 are plotted as a function of time.
We validate the accuracy of our thermal RC and DSS models by comparing their temperature estimates to full-system FEM simulation results for the same workload and system configurations. This comparison is completed for each of the three 2.5D system sizes and the single 3D system. In addition to the visualization of the temperature estimate over time, shown in Fig. 9, three metrics are used to quantify the accuracy of our thermal RC and DSS models against the FEM results. The MAE metric measures the mean absolute error in temperature across the entire simulation duration. The second metric, average percentage error, expresses the temperature error as a percentage. This is particularly important because a given absolute error at lower temperatures represents a more significant inaccuracy than the same absolute error at higher temperatures. The third metric, Temperature Violation Prediction accuracy, measures the accuracy of our models in predicting temperature violations. Predicting temperature violations (e.g., tracking the time steps when the temperature exceeds the allowed threshold) is often used by DTPM algorithms. We set 85C as the maximum allowable temperature threshold for each system without loss of generality (Zhou et al., 2020). This metric first identifies the time steps in FEM simulations where temperature violations occur (temperature exceeds 85C). Then, it computes the percentage of these violations captured by the thermal RC and DSS models (e.g., 100% means all violations are detected with perfect accuracy). The proposed models conservatively flag violations within one degree of the above-mentioned threshold temperature.
To assist users in visualizing the thermal behavior of the system under test, the RC model also creates a heat map of each layer in the system. As an example, the heat map of the interposer layer of a 2.5D 64 chiplet system is shown in Fig. 10. This figure shows the temperature gradient that occurs between the center of the interposer, where the heat-producing chiplets are located, and the edges of the system, where there are no chiplets. These maps allow for quick visual verification of the system behavior instead of relying purely on numerical results.
Figure 10. The heat map of (a) the interposer layer and (b) chiplets generated by the thermal RC model for 64 chiplet system.
2.5D Validation Results: Figs. 9(a), 9(b), and 9(c) plot the temperature as a function of time for each 2.5D system size of a representative chiplet while running workload WL1. All three plots of 2.5D systems clearly show that the systematically constructed thermal RC and DSS models produce near-identical results to the FEM baseline. They closely follow the FEM results during the stress test (temperature increases until reaching the maximum point), randomly changing chiplet power consumption (middle portion), and cool-down periods.
[…]
@@
Accuracy
WL1 Thermal RC 1.23C 1.59% 93.5% 1.42C 1.62% 95.7% 0.99C 1.26% 99.8% 0.94C 1.11% 98.1%
DSS 1.23C 1.59% 93.5% 1.42C 1.62% 95.7% 0.99C 1.26% 99.8% 0.94C 1.11% 98.1%
HotSpot 2.74C 3.78% 67.8% 1.64C 2.13% 95.2% 2.16C 2.62% 95.7% 2.37C 2.99% 87.5%
3D-ICE 1.65C 2.28% 89.5% 2.07C 2.63% 97.4% 1.65C 2.07% 99.5% 1.73C 2.21% 96.9%
PACT 1.99C 2.76% 91.7% 1.61C 2.13% 99.1% 1.70C 2.15% 99.4% 2.09C 2.62% 98.7%
WL2 Thermal RC 0.86C 1.38% 96.0% 1.16C 1.89% 0.0% 0.95C 1.56% 100.0% 0.98C 1.58% 100.0%
DSS 0.86C 1.38% 96.0% 1.16C 1.89% 0.0% 0.95C 1.56% 100.0% 0.98C 1.58% 100.0%
HotSpot 1.58C 2.80% 75.0% 1.24C 1.95% 0.0% 1.01C 1.62% 100.0% 1.74C 2.70% 100.0%
3D-ICE 1.36C 2.21% 85.1% 1.35C 2.13% 0.0% 1.82C 2.93% 100.0% 1.39C 2.21% 100.0%
PACT 1.58C 2.38% 75.9% 0.91C 1.44% 16.7% 1.55C 2.51% 100.0% 2.10C 3.33% 100.0%
WL3 Thermal RC 1.02C 1.63% 100.0% 1.28C 1.93% 78.5% 1.03C 1.57% 95.8% 0.80C 1.23% 100.0%
DSS 1.02C 1.63% 100.0% 1.28C 1.93% 78.5% 1.03C 1.57% 95.8% 0.80C 1.23% 100.0%
HotSpot 1.97C 3.02% 100.0% 1.18C 1.74% 29.7% 1.12C 1.65% 67.4% 1.79C 2.67% 100.0%
3D-ICE 1.15C 1.79% 100.0% 1.53C 2.27% 3.8% 1.82C 2.75% 76.4% 1.12C 1.72% 100.0%
PACT 2.03C 3.12% 100.0% 1.09C 1.61% 23.4% 1.75C 2.59% 66.7% 1.80C 2.76% 100.0%
WL4 Thermal RC 1.46C 1.95% 96.6% 1.64C 2.10% 96.0% 1.35C 1.74% 98.0% 1.48C 1.85% 98.0%
DSS 1.46C 1.95% 96.6% 1.64C 2.10% 96.0% 1.35C 1.74% 98.0% 1.48C 1.85% 98.0%
HotSpot 2.29C 3.36% 95.6% 7.39C 10.28% 92.4% 2.10C 2.59% 96.5% 2.71C 3.36% 96.2%
3D-ICE 3.72C 4.72% 91.3% 2.33C 2.94% 97.2% 2.68C 3.43% 97.3% 2.46C 3.08% 96.4%
PACT 3.56C 4.74% 93.8% 1.94C 2.46% 97.6% 2.39C 3.06% 97.2% 2.97C 3.79% 96.3%
WL5 Thermal RC 1.01C 1.62% 100.0% 1.25C 1.89% 78.4% 0.96C 1.43% 100.0% 0.87C 1.35% 100.0%
DSS 1.01C 1.62% 100.0% 1.25C 1.89% 78.4% 0.96C 1.43% 100.0% 0.87C 1.35% 100.0%
HotSpot 1.95C 3.03% 100.0% 1.16C 1.71% 57.9% 1.30C 1.88% 0.0% 1.86C 2.77% 100.0%
3D-ICE 1.16C 1.82% 100.0% 1.53C 2.27% 15.8% 1.79C 2.67% 50.0% 0.87C 1.33% 100.0%
PACT 1.99C 3.09% 100.0% 1.13C 1.68% 38.4% 1.87C 2.72% 15.0% 1.64C 2.51% 100.0%
WL6 Thermal RC 0.89C 1.44% 98.1% 1.30C 1.97% 84.3% 1.21C 1.79% 89.7% 1.52C 2.36% 91.2%
DSS 0.89C 1.44% 98.1% 1.30C 1.97% 84.3% 1.21C 1.79% 89.7% 1.52C 2.36% 91.2%
HotSpot 1.62C 2.82% 85.9% 1.28C 1.88% 87.8% 1.48C 2.18% 76.1% 2.22C 3.37% 82.3%
3D-ICE 1.34C 2.17% 91.8% 1.69C 2.48% 74.7% 2.38C 3.54% 69.8% 1.61C 2.44% 70.3%
PACT 1.67C 2.56% 89.3% 1.28C 1.91% 91.4% 1.85C 2.75% 90.2% 2.43C 3.69% 66.9%
The first 3 columns of Table 9 summarize the accuracy results for all 2.5D systems and workloads. . The worst-case mean absolute errors of our models are only 1.46C, 1.64C, and 1.35C for the 16-, 36-, and 64-chiplet systems, respectively. The corresponding worst-case average percentage errors are 1.95%, 2.10%, and 1.79%, indicating that even the largest errors occur at higher temperatures where such deviations are less critical. In contrast, other tools exhibit significantly higher errors, as discussed later. These results indicate that the proposed models achieve excellent accuracy across different hardware configurations and workloads. Our models also achieve FEM-level high accuracy in predicting temperature violations as described earlier in Section 5.4. For example, the worst-case accuracy for the 16-chiplet system is 93.5% (i.e., only 6.5% of the time steps where violations are missed). With the exception of WL2 and 3, prediction accuracy exceeds 82% for all other workloads run on the 36 chiplet system. The same trend is seen with the 64-chiplet system.
@@
For some specific workload-system combinations, such as WL2 and 3 run on the 36 chiplet system, prediction accuracy falls below the 80% mark for our models. These relatively lower accuracy values stem from sudden temperature spikes that lead to short-term temperature violations in these workload-system combinations. Temperature spikes in specific chiplets occur when several grouped chiplets experience a transient power spike simultaneously. This increases the temperature of the more central chiplets in the group. In these cases, the peak temperatures are mostly at or below 85 degrees with infrequent short-term violations, which can be tolerated. For example, FEM simulations indicate only 255 temperature violations across all chiplets in WL3 compared to 11 thousand violations in WL1. Hence, missing even a few violations impacts the accuracy heavily when the RC and DSS models do not capture these spikes. In contrast, our thermal RC and DSS models effectively detect more prolonged violations, as evidenced by WL1, WL3, WL4, WL5, and WL6.
@@
3D Validation Results: Fig. 9(d) plots the temperature as a function of time for the single 3D system of a representative chiplet running WL1. The plot shows similar behavior to the 2.5D comparison, where the RC and DSS match each other exactly and match the reference FEM results extremely closely during the stress test, random, and cool-down portions of the workload.
@@
The far-right column of Table 9 summarizes the accuracy results for the single 3D system for each workload. The worst-case mean absolute error is only 1.52C, while the worst-case average percentage error is 2.36%. This result shows that the proposed models maintain high levels of accuracy even when applied to a stacked-die 3D package. This high degree of accuracy is repeated when predicting temperature violations. The worst-case temperature violation prediction accuracy is only 94.0%, occurring for workload 6.
Accuracy comparison to existing tools: Table 9 also lists the MAE of HotSpot, 3D-ICE, and PACT simulations compared to FEM simulations for the 2.5D and 3D systems. For the 2.5D system results, the average error of Hotspot, 3D-ICE, and PACT across all workloads are 0.7, 0.66, and 0.59 degrees greater than our thermal RC and DSS models. For the 3D system results, the average error of Hotspot, 3D-ICE, and PACT across all workloads are 1.02, 0.43, and 1.07 degrees greater than our thermal RC and DSS models.
There are several reasons for the improved accuracy of our models as compared to existing thermal modeling tools. First, PACT and 3D-ICE are not able to model convective heat transfer from more than one boundary of the package. In the case of 3D-ICE, this configuration would require a uniform grid node distribution, which would negatively impact execution time. Second, none of these existing simulators support the ability to provide different thermal conductivities along the x, y, and z-axis, further decreasing accuracy. Third, our models have their node capacitance values tuned, improving their transient thermal simulation accuracy. Lastly, the strength of the PACT simulator lies in circuit-level thermal simulations (Yuan et al., 2021). It assumes a much finer grid granularity compared to block granularity, meaning that each block contains many thermal nodes. However, in architectural-level thermal simulations, the focus of this work, each block contains significantly fewer nodes. This mismatch can lead to uneven node distribution within the PACT simulator, potentially reducing simulation accuracy.
5.5. Modeling of Heterogeneous Architecture
To demonstrate the applicability of MFIT to complex heterogeneous architectures, we model AMD’s MI300A system in this section. The MI300A features a hybrid 2.5D/3D integration, where multiple chiplets are stacked across two tiers (Smith et al., 2024). Specifically, IO dies (IODs) are placed on the bottom tier, with a mix of accelerator complex dies (XCDs) and CPU complex dies (CCDs) stacked directly above them. Our modeled configuration includes 6 XCDs and 3 CCDs. This compute structure is surrounded by 8 high-bandwidth memory (HBM) stacks and their corresponding controllers, as illustrated in Fig. 11(a). We model this system using our thermal RC framework with a total of 500 nodes. As in Section 5.1, we assign higher node density to active components such as the XCDs and CCDs, and lower node density to passive components like the lid and substrate. This demonstrates MFIT’s capability to model complex 2.5D/3D structures with heterogeneous node densities.
Conventional thermal simulators such as HotSpot, PACT, and 3D-ICE are not capable of accurately modeling such architectures. These tools typically assume homogeneous material layers and cannot model anisotropic materials or bidirectional heat dissipation, these capabilities that are critical for capturing the thermal behavior of high-performance industrial systems like MI300A. Fig. 11(b-d) shows temperature heatmaps across three vertical stacks, with labeled components, highlighting the spatial thermal variations across the heterogeneous chiplet layout.
Figure 11. Thermal modeling of AMD MI300A, a CPU-XPU heterogeneous system (a) Block diagram of MI300A architecture (Smith et al., 2024), (b) Heatmap of IO Die layer, (c) Heatmap of compute die layers, (d) Heatmap of bulk-si layer.
6. Conclusion